# Zhang, Liu, and Zhao

### v3

@@ 2.3 Consistency Comparison with Existing DR-based Methods
In comparison, our PADR-based ERM model is nonasymptotically consistent for general SP problems, including nonconvex ones. Moreover, if there exists a convex or DC optimal DR, based on the approximation bounds $\mathcal{O}(K^{-2/p})$ of ${\mathcal{H}}_{\mu}^{K}$ from Balazs et al. (2015) and Siahkamari et al. (2020), our excess risk bound can be improved to $\widetilde{\mathcal{O}}(n^{-1/(3+p/2)})$ with $K=\mathcal{O}(n^{p/(p+6)})$, which is close to the bound of NV-ERM1 (a special case of our model) when $p$ is large. This gap will vanish, that is, the bound can be further refined to $\widetilde{\mathcal{O}}(n^{-1/(2+p/2)})$ for SP with a box-constrained set $[-M,M]^{d}$, by employing a truncated PADR class ¯${\bar{\mathcal{H}}}_{\mu}^{K}\coloneqq\{\bar{f}\mid\bar{f}=\max\{\min\{f,M\},-M\},\,f\in{\mathcal{H}}_{\mu}^{K}\}$ in (3), under which $M_{\mu}$ in Theorem 2.3 could be independent of $\mu$.
Though obtaining the global optimum $f^{*}_{S,{\mathcal{H}}_{\mu}^{K}}$ of the nonconvex ERM problem (5) is prohibitively challenging, we show in Sections 3 and 4 how to asymptotically compute a stationary solution that could be globally optimal to problem (5) under particular algorithmic settings, upon which our PADR-based ERM framework is asymptotically optimal.
3 Enhanced SMM Algorithm for Learning PADR
In this section, we focus on learning PADR by solving problem (5). Following the majorization-minimization (MM) framework, we develop a stochastic algorithm that iteratively solves convex subproblems constructed with sampling-based enhanced surrogation functions, referred to as an enhanced SMM algorithm. The highlight of our algorithm that distinguishes itself from the deterministic MM-related algorithms in Pang et al. (2016) and Cui et al. (2018) and stochastic ones in Xu et al. (2019), Le Thi et al. (2022), and Liu et al. (2022) is that it combines the enhanced $\varepsilon$-active surrogation technique with sequential sampling so as to converge to d-stationary solutions with a significant improvement of computational efficiency.
@@ 3 Enhanced SMM Algorithm for Learning PADR
We first make some remarks on notations to facilitate algorithmic development and convergence analysis. With a slight abuse of notation, to emphasize the role of PADR as a function of $\theta$, we represent the PA function as $f(\theta;\xi)=g(\theta;\xi)-h(\theta;\xi)$ with $\xi\in S=\{(x^{s},y^{s})\}_{s=1}^{n}$, where $g(\theta;\xi)\coloneqq\max_{i\in[K]}g_{i}(\theta;\xi)$ and $h(\theta;\xi)\coloneqq\max_{i\in[K]}h_{i}(\theta;\xi)$ are maximum of linear functions. With $\widetilde{\xi}_{n}$ representing a random variable uniformly distributed on $S$, the PADR-based ERM problem (5) can be generalized as the following stochastic program with finite support:
$\min_{\theta\in\Theta}\quad F(\theta)\coloneqq\mathbb{E}\,\left[\varphi(f(\theta;\widetilde{\xi}_{n});\widetilde{\xi}_{n})\right],$ (9)
@@
where $\Theta\subseteq\mathbb{R}^{q}$ is a compact convex set. In particular, in problem (5) with $\Theta=[-\mu,\mu]^{q}$ and $q=2dK(p+1)$, the outer function $\varphi(\cdot;\xi)$ with $\xi=(x,y)$ depends solely on $y$, and the inner PA function $f(\cdot;\xi)$ depends solely on $x$.
3.1 Construction of Surrogation Functions
Similar to other MM-related algorithms, the cornerstone of the enhanced SMM algorithm is the construction of surrogation functions of $F(\cdot;\xi)\coloneqq\varphi(f(\cdot;\xi);\xi)$. For a directionally differentiable function $\phi:\Theta\to\mathbb{R}$, that is, the directional derivative $\phi^{\prime}(\theta;d)\coloneqq\lim_{\tau\to 0}\frac{\phi(\theta+\tau d)-\phi(\theta)}{\tau}$ exists for any $\theta\in\Theta$ and $d\in\mathbb{R}^{q}$, we are interested in constructing a surrogation function $\widehat{\phi}(\cdot;\theta^{\prime})$ at a reference point $\theta^{\prime}\in\Theta$ that may satisfy the following conditions:
@@ 3.1 Construction of Surrogation Functions
    Touching condition: $\widehat{\phi}(\theta^{\prime};\theta^{\prime})=\phi(\theta^{\prime})$.
@@
    Majorizing condition: $\widehat{\phi}({\theta};{\theta^{\prime}})\geq\phi(\theta)$ for any $\theta\in\Theta$.
@@
    Convexity: $\widehat{\phi}({\cdot};{\theta^{\prime}})$ is convex on $\Theta$.
  4. (P4)
    Directional derivative consistency: $\widehat{\phi}(\cdot;\theta^{\prime})$ is directionally differentiable on $\Theta$ and satisfies $\widehat{\phi}(\cdot;\theta^{\prime})^{\prime}(\theta^{\prime};d)=\phi^{\prime}(\theta^{\prime};d)$ for any $d\in\mathbb{R}^{q}$.
  5. (P5)
    Quadratic surrogation gap: for some $\kappa>0$, $\phi(\theta)+\displaystyle{\frac{\kappa}{2}}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}\geq\widehat{\phi}(\theta;\theta^{\prime})$ for any $\theta\in\Theta$.
A function satisfying the first two conditions is referred to as an upper surrogation function according to Cui and Pang (2021, Definition 7.1.1). Though the enhanced $\varepsilon$-active surrogation functions to be constructed generally do not satisfy (P1), we still refer to ones solely satisfying (P2) as upper surrogation functions and similarly for lower surrogation functions. Condition (P3) ensures that surrogation-based subproblems in the algorithm can be solved by convex-programming solvers. Condition (P4) connecting the directional derivatives of surrogation and original functions is used to establish stationarity conditions of the limit points produced by surrogation-based algorithms to the original problem. Condition (P5) originating from Liu et al. (2022) is a technical condition for the nonasymptotic convergence analysis of the enhanced SMM algorithm. Note that Condition (P5) implies (P4) provided that $\widehat{\phi}(\cdot;\theta^{\prime})$ is directionally differentiable and satisfies (P2).
We first present upper and lower surrogation functions of the inner PA function $f(\cdot;\xi)$, which provide the basis of surrogation construction of $F(\cdot;\xi)$ and also play an essential role in the randomized index-mapping selection in the enhanced SMM algorithm. For $d=1$ and $\varepsilon\geq 0$, let $\mathcal{I}_{g}^{\varepsilon}(\theta^{\prime})$ denote the set of $\varepsilon$-active index mappings for functions $\{g(\cdot;\xi)\}_{\xi\in S}$ at $\theta^{\prime}\in\Theta$, i.e.,
$\mathcal{I}_{g}^{\varepsilon}(\theta^{\prime})\coloneqq\Big{\{}I_{1}:S\rightarrow[K]\mid\,g_{I_{1}(\xi)}(\theta^{\prime};\xi)\geq g(\theta^{\prime};\xi)-\varepsilon,\,\forall\,\xi\in S\Big{\}},$
and the set $\mathcal{I}_{h}^{\varepsilon}(\theta^{\prime})$ is similarly defined for $\{h(\cdot;\xi)\}_{\xi\in S}$. Each mapping $I_{1}\in\mathcal{I}_{g}^{\varepsilon}(\theta^{\prime})$ assigns an $\varepsilon$-active index for each $\xi\in S$, and thus $\{I_{1}(\xi)\}_{\xi\in S}$ constitutes an $\varepsilon$-active index combination of $\{g(\cdot;\xi)\}_{\xi\in S}$ at $\theta^{\prime}$. It is clear to see that $\widehat{f}(\theta;\xi,I_{2})\coloneqq g(\theta;\xi)-h_{I_{2}(\xi)}(\theta;\xi)$ with $I_{2}\in\mathcal{I}_{h}^{\varepsilon}(\theta^{\prime})$ is an upper surrogation function of $f(\cdot;\xi)$ satisfying Conditions (P2) and (P3) with $\varepsilon\geq 0$, and additionally (P1) with $\varepsilon=0$. The concave lower surrogation function $\widecheck{f}(\theta;\xi,I_{1})$ with $I_{1}\in\mathcal{I}_{g}^{\varepsilon}(\theta^{\prime})$ can be similarly defined. When $d>1$, the index mapping sets and surrogates are constructed componentwisely.
Now, we can construct the upper surrogation function of $F(\cdot;\xi)$ at a reference point $\theta^{\prime}$ based on the surrogation of inner PA functions with the index mapping $I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})\coloneqq\mathcal{I}_{g}^{\varepsilon}(\theta^{\prime})\times\mathcal{I}_{h}^{\varepsilon}(\theta^{\prime})$, which we denote by $\widehat{F}(\cdot;\theta^{\prime},\xi,I)$. To facilitate the convergence analysis of the enhanced SMM algorithm, we impose the following assumptions on the family of surrogates $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$:
  1. [label=(B0), series=B]
  2. 1.
    The family of surrogates $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$ for any $\xi\in S$ satisfy the following properties:
    1. (a)

### v4

@@ 2.3 Consistency comparison with existing DR-based methods
In comparison, our PADR-based ERM model is nonasymptotically consistent in more general cases, including convex SP problems and nonconvex ones. Moreover, by Corollary 2.4, if there exists a convex or DC optimal DR, our excess risk bound can be improved to $\widetilde{\mathcal{O}}(n^{-1/(3+p/2)})$, which is close to the bound of NV-ERM1 (a special case of our model) when $p$ is large. This gap will vanish, that is, the bound can be further refined to $\widetilde{\mathcal{O}}(n^{-1/(2+p/2)})$, for problem (1) with a box-constrained set $[-M,M]^{d}$ by employing a truncated PADR class ¯${\bar{\mathcal{H}}}_{\mu}^{K}\coloneqq\{\bar{f}\mid\bar{f}=\max\{\min\{f,M\},-M\},\,f\in{\mathcal{H}}_{\mu}^{K}\}$ in (3), under which $M_{\mu}$ in Theorem 2.3 could be independent of $\mu$.
Given the statistical consistency merits of the PADR method demonstrated by the analysis above, it is still prohibitively challenging to obtain the globally optimal solution to the nonconvex PADR-based ERM problem (5). From the continuous optimization perspective, the PADR-based ERM problem is a composite nonconvex and nondifferentiable optimization problem embedded with a particular piecewise affine structure. This leads to the development of a sampling-based surrogate-based algorithm in the next section to efficiently solve problem (5) of large sizes.
3 Efficient Sampling-based Algorithm for Learning PADR
In this section, we focus on learning PADR by solving the following parameter estimation problem
$\min_{\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu,\forall\iota\in[d]}\quad\frac{1}{n}\sum_{s=1}^{n}\varphi\left(\Big{(}\max_{k\in[K]}\left\{(\alpha^{k}_{\iota})^{\top}x^{s}+a^{k}_{\iota}\right\}-\max_{k\in[K]}\left\{(\beta^{k}_{\iota})^{\top}x^{s}+b^{k}_{\iota}\right\}\Big{)}_{\iota=1}^{d},y^{s}\right),$ (9)
where the objective is the finite sum of composite-PA functions. Instead of attempting for global optimality, it is more practical to apply majorization-minimization (MM)-type methods to the contextual optimization problems with the practical scale in seeking “stationary” solutions. Among various kinds of stationary concepts, the d-stationary solution is the sharpest type of first-order stationary solutions for general nonconvex optimization. For the PADR learning problem (9) with a general nonconvex and nondifferentiable composite structure, our algorithmic development is mainly motivated by the following question:
Given the potentially large sample size, how to efficiently compute a sharper “stationary” solution to the PADR-based ERM problem (9) so that a global minimizer could be more easily obtained through replications with random initialization?
Building on this motivation, we develop an MM-type algorithm specializing to the structured PADR-based ERM problem with the theoretical convergence analysis. It is further observed with numerous experiments in Section 4 that satisfactory (nearly optimal) solutions can be obtained with a few number of repetitions, which speeds up to 4 times faster than the basic MM method.
@@ 3 Efficient Sampling-based Algorithm for Learning PADR
To make our discussion self-contained, we first present the standard MM algorithms in Section 3.1, which serves as a reference for the motivation behind our proposed sampling-based algorithm in Section 3.2. We make some remarks on notations to facilitate algorithmic development and convergence analysis. To emphasize the role of PADR as a function of $\theta$, with $\xi=(\xi_{x},\xi_{y})\in\Xi$, we denote $g(\theta,\xi)\coloneqq\max_{i\in[K]}g_{i}(\theta,\xi_{x})$ and $h(\theta,\xi)\coloneqq\max_{i\in[K]}h_{i}(\theta,\xi_{x})$ where $\{g_{i}(\theta,\xi_{x})\}$ and $\{h_{i}(\theta,\xi_{x})\}$ are linear functions with respect to $\theta$ and we denote $f(\theta,\xi)\coloneqq g(\theta,\xi)-h(\theta,\xi)$ as the PA function with a slight abuse of notation.
With $\widetilde{\xi}_{n}$ representing a random variable uniformly distributed on $\Xi$, the PADR-based ERM problem (5) can be generalized as the following stochastic program with finite scenarios:
$\min_{\theta\in\Theta}\quad F(\theta)\coloneqq\mathbb{E}\,\left[F(\theta,\widetilde{\xi}_{n})\right],$ (10)
@@
where $F(\theta,\xi)\coloneqq\varphi(f(\theta,{\xi}),{\xi})$ and $\Theta\subseteq\mathbb{R}^{q}$ is a compact convex set. In particular, it reduces to problem (5) when $\Theta=[-\mu,\mu]^{q}$ with $q=2dK(p+1)$, the outer function $\varphi(\cdot,\xi)$ with $\xi=(x,y)$ depends solely on $y$, and the inner PA function $f(\cdot,\xi)$ depends solely on $x$. We say a function $F:\Theta\subseteq\mathbb{R}^{q}\to\mathbb{R}$ is directionally differentiable if the directional derivative $F^{\prime}(\theta;d)\coloneqq\lim_{\tau\to 0}\frac{F(\theta+\tau d)-F(\theta)}{\tau}$ exists for any $\theta\in\Theta$ and $d\in\mathbb{R}^{q}$. Supposing that $\Theta$ is a closed convex set and $F:\Theta\rightarrow\mathbb{R}$ is directionally differentiable, we say that $\theta^{d}\in\Theta$ is a d(irectional)-stationary point of problem (10) if $F^{\prime}(\theta^{d};\theta-\theta^{d})\geq 0$ for any $\theta\in\Theta$.
3.1 Preliminary analysis: standard MM algorithms
The core of MM-type methods for solving nonconvex minimization problem is to iteratively solve subproblems with a surrogate function $\widehat{F}(\theta;\theta^{\nu})$ majorizing $F(\theta)$ at an iterate point $\theta^{\nu}\in\Theta$. The iterate point is updated with $\eta\geq 0$ as follows,
$\theta^{\nu+1}\leftarrow\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}(\theta;\theta^{\nu})+\frac{\eta}{2}\left\lVert\theta-\theta^{\nu}\right\rVert^{2}\right\}.$ (11)
A surrogate function $\widehat{F}(\theta;\bar{\theta})$ at a reference point $\bar{\theta}\in\Theta$ typically satisfies the following conditions:
@@ 3.1 Preliminary analysis: standard MM algorithms
    Touching condition: $\widehat{F}(\bar{\theta};\bar{\theta})=F(\bar{\theta})$.
@@
    Majorizing condition: $\widehat{F}({\theta};{\bar{\theta}})\geq F(\theta)$ for any $\theta\in\Theta$.
@@
    Convexity: $\widehat{F}({\cdot};{\bar{\theta}})$ is convex on $\Theta$.
A function satisfying (P1) and (P2) is referred to as an upper surrogate function according to Cui and Pang (2021, Definition 7.1.1). Under (P1) and (P2), it yields the sequence of decreasing objective values as follows,
$F(\theta^{\nu+1})\leq\widehat{F}(\theta^{\nu+1};\theta^{\nu})\leq\widehat{F}(\theta^{\nu};\theta^{\nu})-\frac{\eta}{2}\left\lVert\theta^{\nu+1}-\theta^{\nu}\right\rVert^{2}=F(\theta^{\nu})-\frac{\eta}{2}\left\lVert\theta^{\nu+1}-\theta^{\nu}\right\rVert^{2}\leq F(\theta^{\nu}).$
Condition (P3) ensures that surrogation-based subproblems in the algorithm can be solved by state-of-the-art convex-programming solvers. By the basic analysis (see Chapter 7 in Cui and Pang (2021)), the limit point of $\{\theta^{\nu}\}$ produced by the process (11) could be easily shown to be a fixed point of the surrogation map satisfying $\theta^{*}=\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}(\theta;\theta^{*})+\frac{\eta}{2}\left\lVert\theta-\theta^{*}\right\rVert^{2}\right\}$. Such a fixed point property could be related to various types of stationarity of the original problem depending on the surrogate functions as illustrated in Cui et al. (2018) and Liu et al. (2022).
Regarding the PADR-based ERM problem (10), due to the nonconvex and nondifferentiable piecewise affine functions, a major issue of the standard MM algorithms including Lipp and Boyd (2016) and Le Thi and Pham Dinh (2018) is that the fixed point corresponding to the surrogation map is a (compound) critical solution (see definitions in Pang et al. (2016), Liu et al. (2022)) which is much weaker than directional stationarity. Specifically, based on the composite structure in (10), approaches to construct surrogate functions of $F(\theta,\xi)$ in Cui et al. (2018) typically involve linearization of the two max-affine components of the inner piecewise affine function as follows,
$\widehat{f}(\theta,\xi;I_{2})$ $\coloneqq g(\theta,\xi)-h_{I_{2}}(\theta,\xi)\mbox{ with }I_{2}\in\mathcal{I}_{h}(\bar{\theta},\xi)\coloneqq\mbox{argmax}_{i\in[K]}h_{i}(\bar{\theta},\xi),$ (12)
$\widecheck{f}(\theta,\xi;I_{1})$ $\coloneqq g_{I_{1}}(\theta,\xi)-h(\theta,\xi)\mbox{ with }I_{1}\in\mathcal{I}_{g}(\bar{\theta},\xi)\coloneqq\mbox{argmax}_{i\in[K]}g_{i}(\bar{\theta},\xi).$ (13)
They are the upper and lower surrogate functions of $f(\cdot,\xi)$ at $\bar{\theta}\in\Theta$ respectively. For $d>1$, the maximum is taken componentwisely. We illustrate the construction of the surrogate function when outer function $\varphi$ is linear as an example. With $\varphi(z,\xi)=\xi^{\top}z$, we construct the surrogate function of $F(\cdot,\xi)$ as $\widehat{F}(\theta,\xi;\bar{\theta},I)=\max\{\xi,0\}^{\top}\widehat{f}(\theta,\xi;I_{2})+\min\{\xi,0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})$, where $I\coloneqq\big{(}I_{1},I_{2}\big{)}\in\mathcal{I}_{g}(\bar{\theta},\xi)\times\mathcal{I}_{h}(\bar{\theta},\xi)$ for any $\xi\in\Xi$. It obviously satisfies Conditions (P1) to (P3), and the standard MM algorithm for such a case is exactly a DC algorithm. Clearly $f^{\prime}(\cdot,\xi)(\theta;d)=g^{\prime}(\cdot,\xi)(\theta;d)-\max\{\,\nabla h_{i}(\bar{\theta},\xi)^{\top}d:\,i\in\mbox{argmax}_{i\in[K]}h_{i}(\bar{\theta},\xi)\}$, whereas the surrogate function $\widehat{F}(\theta,\xi;\bar{\theta},I)$ employs an arbitrary active index $I\in\mathcal{I}_{g}(\bar{\theta},\xi)\times\mathcal{I}_{h}(\bar{\theta},\xi)$. Hence, the fixed point of the surrogation map is a critical solution. Solutions of criticality are not desirable because in practice solutions may be at the “cliff” of the cost landscape, which is illustrated in Example 3.1 and Figure 1b.
To obtain a directional stationary solution, Pang et al. (2016) considered a class of nondifferentiable DC programs which have a max-smooth structure in the second convex part, and introduced $\varepsilon$-active sets to construct the enhanced surrogate functions for DC functions. Specifically, for the piecewise affine function $f(\cdot,\xi)$ given $\xi\in\Xi$, we denote the $\varepsilon$-active set with $\varepsilon>0$ by
$\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\coloneqq\Big{\{}i\in[K]\mid\,g_{i}(\bar{\theta},\xi)\geq g(\bar{\theta},\xi)-\varepsilon\Big{\}},$ (14)
and the set $\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)$ is similarly defined for $h(\cdot,\xi)$. Let $\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)\coloneqq\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)$. Pang et al. (2016) proposes the enhanced MM algorithm in which the subproblem takes the minimum over the family of surrogate functions with indices in $\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)$ and provides the theoretical convergence analysis to a d-stationarity point. The following example provides an intuitive illustration that the iterate point can escape from a critical point to better local minima and make valid descents via the enhanced technique.
Example 3.1
Consider problem (10) with $d=2$, $q=1$, $\Xi=\{\xi_{0}\}$, $\varphi(z,\xi_{0})=z_{1}^{2}+z_{2},f_{1}(\theta,\xi_{0})=\theta$, and $f_{2}(\theta,\xi_{0})=-\max\{-\theta,0\}$. Figure 1a illustrates the two candidates of surrogate functions depending on which piece of the inner PA function $f_{2}(\theta,\xi_{0})$ is active at the reference point of interest. Figure 1b and Figure 1c illustrate the iterative procedures of the standard MM algorithm and the enhanced MM algorithm respectively, with the initial point $\theta_{0}=1.1$. The standard MM algorithm converges to a critical point $\theta_{\infty}=0$. For the enhanced MM algorithm with $\varepsilon=0.3$, at the second iterate point $\theta_{2}$, we have $\left\lvert\mathcal{I}^{\varepsilon}(\theta_{2},\xi_{0})\right\rvert=2$ at $\theta_{2}$, which thus leads to the next iterate point passing over 0 and continues to move towards a d-stationary point $\theta^{d}$ that is also globally optimal to the problem.
[…]
(a) DC function
(b) Surrogation with $\varepsilon=0$
(c) Surrogation with $\varepsilon=0.3$
Iterative process of the standard MM algorithm and enhanced MM algorithm
Despite possessing stronger convergence properties, the enhanced MM algorithm applied for large-scale ERM problems could be exceptionally inefficient as it requires iteratively solving a potentially very large number of finite-sum convex subproblems over all the data. Sampling-based MM-type algorithms have been recently studied for (partially) smooth stochastic DC programs in Nitanda and Suzuki (2017), Xu et al. (2019), and Le Thi et al. (2022) and nondifferentiable stochastic composite DC problems in Liu et al. (2022) to compute compound critical solutions. Based on a natural idea of combining a sequential sampling scheme with the enhanced technique, we propose an enhanced sampling-based MM algorithm with rigorous convergence analysis in the next subsection to solve the stochastic composite DC programs (10) with both computational efficiency and theoretical convergence guarantee.
3.2 Enhanced sampling-based MM algorithm for learning PADR
In this subsection, we first introduce the family of enhanced surrogate functions of the composite objective $F$ and provide several function classes of the outer function $\varphi$ as particular instances. This provides the basis for the characterization of a so-called composite $(\varepsilon,\rho)$-strong d-stationarity. We then propose the enhanced sampling-based MM algorithm (ESMM) which combines the enhanced $\varepsilon$-active technique with sequential sampling and analyze the asymptotic and non-asymptotic convergence properties in terms of (composite strong) d-stationarity under suitable incremental sampling schemes. Such a combination enables the algorithm to solve subproblems of small sizes in the early iterations that effectively reduces computational costs. The computational improvement of such a technique is further supported by numerical experiments in Section 4.
3.2.1 Family of enhanced surrogate functions.
Based on surrogation of $f(\theta,\xi)$ in (12) and (13) and corresponding $\varepsilon$-active index set $\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)\coloneqq\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)$ given $\bar{\theta}\in\Theta$, we introduce a family of enhanced surrogate functions $\{\widehat{F}(\cdot,\xi;\bar{\theta},I)\}_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\,\xi)}$ for any $\xi\in\Xi$ that satisfy the following conditions:
  1. (B1)
    $\widehat{F}(\cdot,\xi;\bar{\theta},I)$ satisfies (P1) for every $I\in\mathcal{I}^{0}(\bar{\theta},\,\xi)$;
  2. (B2)